[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu May 11, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu May 11 07:27:34 UTC 2017


HSU Shift Log for Thu May 11 2017 03:27:28 GMT-0400 (EDT)

Shift: A

Forecaster: L Avila

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Adrian most likely dissipate this morning if is not  
already dissipated.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170511/cde00a87/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list